# Geometry-aware Temporal Aggregation Network for Monocular 3D Lane Detection

### v1

@@ Abstract
Monocular 3D lane detection aims to estimate 3D position of lanes from frontal-view (FV) images. However, current monocular 3D lane detection methods suffer from two limitations, including inaccurate geometric information of the predicted 3D lanes and difficulties in maintaining lane integrity. To address these issues, we seek to fully exploit the potential of multiple input frames. First, we aim at enhancing the ability to perceive the geometry of scenes by leveraging temporal geometric consistency. Second, we strive to improve the integrity of lanes by revealing more instance information from temporal sequences. Therefore, we propose a novel Geometry-aware Temporal Aggregation Network (GTA-Net) for monocular 3D lane detection. On one hand, we develop the Temporal Geometry Enhancement Module (TGEM), which exploits geometric consistency across successive frames, facilitating effective geometry perception. On the other hand, we present the Temporal Instance-aware Query Generation (TIQG), which strategically incorporates temporal cues into query generation, thereby enabling the exploration of comprehensive instance information. Experiments demonstrate that our GTA-Net achieves SoTA results, surpassing existing monocular 3D lane detection solutions.
@@ 1 Introduction
3D lane detection stands as a cornerstone within the realm of autonomous driving systems. Its primary objective involves the meticulous estimation of lane line 3D coordinates [36]. Accurate perception of 3D lanes across diverse environments is crucial, not only for ensuring adherence to lane boundaries and consistent lane keeping [46], but also for enabling important downstream tasks such as high-definition map construction [23, 32, 27] and trajectory planning [50, 1]. Considering the cost-effectiveness of sensors and the rich color information provided by visual data, monocular vision-based 3D lane detection has become a pivotal research direction in autonomous driving [36].
Figure 1: Comparison with Monocular 3D Lane Detection Approaches. (a) Existing methods [17, 2, 13, 5, 10] that rely on a single frame as input are hindered by inaccuracies in 3D geometry perception and difficulties in maintaining lane integrity. (b) In contrast, our approach, which incorporates multiple frames along with synthetic future frames, facilitates improved 3D geometric perception and enables integral lane detection within the scene.
Estimating lanes in 3D space from a 2D image is a formidable challenge, primarily due to the lack of depth cues. To address this, some methods decompose 3D lane detection into two stages: first detecting lanes in 2D, then projecting them into 3D space [54]. Specifically, SALAD [54] tackles this problem by performing two tasks: 2D lane segmentation and dense depth estimation, which together facilitate 3D lane detection. After segmenting the 2D lanes, they are projected into 3D space using camera intrinsics and the predicted depth information. However, this approach still faces two significant drawbacks: it relies heavily on the accuracy of depth estimation and demonstrates suboptimal performance in 3D lane detection.
An alternative approach involves detecting lanes from a bird’s-eye-view (BEV) perspective [5, 10, 13, 31, 48]. These BEV-based methods decompose 3D lane detection into two tasks: 2D lane detection in BEV space and corresponding height estimation. Typically, they employ inverse perspective mapping (IPM) [37] to warp image features from the frontal-view (FV) space to BEV, effectively transforming perspective-view (PV) features into BEV features. To project the detected BEV lanes into 3D space, the BEV coordinates are combined with the corresponding height values which serve as representations of the 3D lanes. However, these BEV-based methods still suffer from the following limitations. Firstly, IPM operates under the assumption of a level surface, making it less effective in scenarios involving uphill or downhill slopes where the road surface deviates from flat terrain. Secondly, due to its reliance on ground-based image warping, IPM inevitably sacrifices valuable height details and contextual information above the road surface. Thirdly, BEV features, unlike perspective features, lack the ability to perceive distant information, thus exhibiting a more limited perception range.
To tackle the aforementioned challenges, recent methodologies incorporate 3D anchors or queries to sample FV features in the context of monocular 3D lane detection [17, 35, 2, 3]. These 3D anchors or queries serve as pivotal elements, used to generate 2D coordinates in image coordinates to sample relevant features. Subsequently, cross-attention mechanisms are employed, allowing for information interaction and iterative refinement of these anchors or queries. This integration is capable of providing a more comprehensive understanding of the spatial context, ultimately leading to improved performance for monocular 3D lane detection in challenging scenarios.
However, current monocular 3D lane detection methods are constrained by two limitations, including the inaccurate geometric representation of predicted 3D lanes and the difficulty in maintaining lane integrity. These challenges may lead to suboptimal performance in real-world applications. As shown in figure 1 (a), the predicted results exhibit the following issues: errors in the positioning of the lanes within 3D space, and incomplete detection, with some lanes being only partially identified or even missed entirely. In this paper, we aim to address these limitations by fully leveraging the potential of multiple input frames from two complementary perspectives, as expressed in figure 1 (b). Firstly, we focus on enhancing the ability to perceive the geometry of scenes by exploiting temporal geometric consistency across consecutive frames. This allows for more accurate and stable estimations of lane positions in 3D space. Secondly, we seek to improve the integrity of lanes by leveraging temporal sequences to extract comprehensive lane instance information. In detail, we leverage images from historical frames and a synthesized future frame to capture instance information for both near and distant regions, thereby mitigating uncertainties in these areas to improve lane integrity.
Based on the aforementioned insight, we propose an end-to-end monocular 3D lane detection framework, named Geometry-aware Temporal Aggregation Network (GTA-Net). To be specific, we introduce two novel modules in our work including Temporal Geometry Enhancement Module (TGEM) and Temporal Instance-aware Query Generation (TIQG). TGEM effectively leverages geometric consistency across consecutive frames, enabling accurate geometric perception of the scene. Meanwhile, TIQG strategically incorporates temporal cues into query generation, facilitating the comprehensive utilization of temporal instance information to enhance lane integrity. These generated queries from TIQG, along with the geometry-aware features from TGEM, are then fed into a deformable attention-based decoder, facilitating iterative information interaction. Finally, a Multilayer Perceptron (MLP)-based 3D lane detection head is employed to produce the final 3D lane predictions.
In brief, the paper presents the following contributions:
@@
  • We propose the Geometry-aware Temporal Aggregation Network (GTA-Net) for monocular 3D lane detection, which achieves highly reliable predictions by effectively leveraging temporal information.
@@
  • We introduce the Temporal Geometry Enhancement Module (TGEM) to enhance the geometric perception of the scene by exploiting the geometric consistency between consecutive frames.
  • We develop the Temporal Instance-aware Query Generation (TIQG) to generate lane queries by fully utilizing the potential of instance information in temporal cues to alleviate difficulties in maintaining lane integrity.
  • Experimental results show that the proposed GTA-Net outperforms existing monocular 3D lane detection methods, achieving state-of-the-art performance.
@@ 2 Related Work
2D lane detection, a critical aspect of autonomous driving, strives to precisely determine the shapes and positions of lanes within images [45, 53]. With the development of deep learning [4, 8, 51, 21], deep neural network-based 2D lane detection has achieved remarkable strides forward [20, 40, 47, 59, 14]. According to the formulation of 2D lane detection, existing solutions can be divided into four categories: segmentation-based, anchor-based, keypoint-based and curve-based methods [17].
Segmentation-based approaches [22, 24, 58] employ segmentation tasks as intermediaries for executing lane detection. To be specific, these methods firstly complete pixel-wise classification to obtain 2D lane segmentation masks. Next, the segmentation masks are processed by some carefully-design strategies to construct a set of lanes. Anchor-based approaches primarily utilize line-like anchors for regressing offsets relative to ground truth lanes [19, 29, 39, 40, 43, 56]. To enhance the efficacy of line-shape anchors, row-based anchors have been devised for capitalizing on the natural shape of lanes [56]. By using row-based anchors, 2D lane detection task is converted into a row-wise classification problem. Furthermore, these row-based and column-based anchors are combined to reduce the occurrence of localization errors in side lanes [40]. Keypoint-based approaches offer a flexible framework for modeling lane positions [22, 41, 47]. At the outset, these methodologies embark on predicting the 2D coordinates of pivotal points along the lanes [22]. Following this initial step, a diverse array of strategies is employed to effectively group these key points into distinct lanes [47]. Curve-based approaches leverage a range of curve formulations to model lanes [9, 30, 44]. Specifically, this type of method aims to estimate the parameters of curves to obtain 2D lane detection results. BézierLaneNet [9] leverages the parametric Bézier curve for its computational simplicity, stability, and extensive flexibility in transformations. Additionally, the authors introduce a novel feature flip fusion module, which takes into account the symmetrical characteristics of lanes within driving environments.
@@
The objective of 3D lane detection is to identify lanes within 3D space and accurately estimate their coordinates [36, 60, 25]. In contrast to 2D lane detection, 3D lane detection methods need to perceive the geometric information of scenes. Some methodologies try to enhance perception accuracy by leveraging both lidar and camera data [34, 33]. However, the acquisition and annotation expenses associated with multi-modal data are prohibitively high, constraining the practical implementation of such methodologies. Hence, monocular 3D lane detection garners increased interest due to its provision of a cost-effective alternative to multi-modal methodologies [36].
Monocular 3D lane detection is challenging due to the absence of depth information. To address this issue, SALAD [54] incorporates a depth prediction branch to estimate depth, which is used to project 2D lanes to 3D space. However, employing depth estimation in the process of detecting lanes may introduce cumulative errors and further exhibits inferior performance. IPM-based approaches have become increasingly prominent owing to their advantageous geometric properties for lane depiction in BEV perspective [36]. 3D-LaneNet [10] utilizes IPM for the conversion of FV features into BEV features, enabling bidirectional information flow between image-view and top-view representations. Next, these features are leveraged for learning the anchor offsets of lanes in the BEV space. Gen-LaneNet [13] introduces a novel lane anchor representation guided by geometry in a newly defined coordinate system, which employs a specific geometric transformation to directly compute real 3D lane points from the network output. It is essential to emphasize that aligning the lane points with the underlying top-view features in the newly defined coordinate frame is crucial for establishing a generalized method capable of effectively handling diverse scenes.
IPM relies heavily on the ground plane assumption, which cannot be guaranteed in real scenarios [6]. To solve this issue, Anchor3DLane [17] is introduced, eliminating the need for BEV representation and instead utilizing 3D lane anchors directly in 3D space. These 3D lane anchors are projected onto image coordinate system for feature extraction, capturing both structural and contextual information crucial for precise lane detection. LATR [35] uses a transformer-based architecture for monocular 3D lane detection. Specifically, the authors design a lane-aware query generator that produces queries derived from 2D lane segmentation. Besides, dynamic 3D ground PE generator is presented to embed 3D information during query update.
Figure 2: Illustration of the Overall Architecture of the Proposed GTA-Net. The input images are first fed into a backbone network to extract 2D perspective features. Next, TGEM is applied to enhance the geometric perception of the scene. Simultaneously, TIQG is used to generate lane queries. These lane queries and features are sent to the lane decoder to obtain the final 3D lane predictions.
@@ 3 Method
In this section, we will initially present a comprehensive introduction to the proposed Geometry-aware Temporal Aggregation Network (GTA-Net) in Section 3.1. To enhance the understanding of the geometric structure inherent in driving scenes, we introduce Temporal Geometry Enhancement Module (TGEM) in Section 3.2. Following this, we introduce Temporal Instance-aware Query Generation (TIQG) aimed at generating queries by fully utilizing temporal information in Section 3.3. Finally, we delve into the employed loss functions in Section 3.4.
@@ 3.1 Overall Architecture
$L=(P,C),\quad P={(x_{i},y_{i},z_{i})}_{i=1}^{N},\$ (1)
where $L$ denotes a 3D lane, $C$ represents the corresponding lane category of $L$, $P$ is composed of $N$ 3D points with $N$ being a predefined parameter. It is noted that the sequence $\{y_{1},y^{2},...,y^{N}\}$ is often defined as a predetermined longitudinal coordinate [10, 13, 5]. Based on the above descriptions about lane representation, the 3D lanes in an input image $I_{t}$ can be represented as $G=\{L^{1},L^{2},...,L^{M}\}$, where $G$ represent the collection of all 3D lanes, and $M$ denotes the number of 3D lanes.
@@
$F_{t}=\text{Backbone}(I_{t}),\quad F_{t-n}=\text{Backbone}(I_{t-n}),\$ (2)
@@
$Q=\text{TIQG}(F_{t-n},F_{t}),\$ (4)
where $Q$ represents the generated instance-aware queries, and $\text{TIQG}(\cdot)$ denotes the transformation performed by the TIQG module. Next, the instance-aware queries $Q$ and geometry-aware features $F_{ge}$ are fed into a deformable attention [52, 61] based decoder to iteratively update queries:
$Q_{u}=\text{Decoder}(Q,F_{ge}),\$ (5)
where $Q_{u}$ and $\text{Decoder}(\cdot)$ denote the updated queries and the transformation of the above deformable attention-based decoder, respectively. Finally, the updated queries $Q_{u}$ are fed into the 3D lane detection head to output the predictions of 3D lanes in the input image $I_{t}$, which can be described as follows:
$\hat{G}=\text{Head}(Q_{u}),\$ (6)
where $\hat{G}$ indicates the estimated 3D lanes, and $\text{Head}(\cdot)$ denotes 3D lane detection head.
@@ 3.2 Temporal Geometry Enhancement Module
The inherent absence of depth cues in monocular input images poses a significant challenge for monocular 3D lane detection solutions. Without depth information, accurately understanding the complex geometric details of driving scenes becomes difficult. Drawing inspiration from prior research [49, 55] in monocular depth estimation, this study aims to alleviate this limitation by utilizing the geometric relationship between consecutive frames. To this end, we propose Temporal Geometry Enhancement Module (TGEM), which is designed to delve into the geometric consistency hidden in successive frames. The structure of TGEM is described by figure 2. In detail, TGEM consists of three main stages: cost volume construction, geometric feature leaning and geometry-aware feature enhancement.
Cost volume construction. To uncover the potential geometric information within successive monocular images, we first construct a cost volume in the initial stage of TGEM. This is because the cost volume is capable of capturing the geometric relationship between adjacent frames [11].
@@
$E_{t}=\text{Repeat}(F_{t}),\quad E_{t-n}=\text{Repeat}(F_{t-n}),\$ (7)
where $E_{t}$ and $E_{t-n}$ denote the expanded features, and $\text{Repeat}(\cdot)$ represents the feature expanding operation. Next, the expanded features $E_{t-n}$ in image coordinate system of $t-n$ frame is projected onto image coordinate system of $t$ frame by using camera intrinsics, extrinsics and lidar poses:
$\hat{E_{t}}=\text{Warp}(F_{t-n}),\$ (8)
where $\hat{E_{t}}$ denotes the warped features in image coordinate system of $t$ frame, and $\text{Warp}(\cdot)$ indicates the projection from image coordinate system of $t-n$ frame to image coordinate system of $t$ frame. In the final step, the matching cost between $\hat{E_{t}}$ and $E_{t}$ is computed by using $l_{1}$ norm:
$C=\lVert\hat{E_{t}}-E_{t}\rVert_{1},\$ (9)
@@
Figure 3: An Overview of the Proposed Temporal Instance-aware Query Generation (TIQG) Module. TIQG strategically integrates temporal cues into the query generation process, thereby enhancing its capacity to exploit temporal instance information for the promotion of lane integrity.
@@
$F_{g}=\text{ReLU}(\text{BN}(\text{Conv}(C))),\$ (10)
@@
$\alpha=\text{Sigmoid}(\text{BN}(\text{Conv}(F_{g}))),\$ (11)
$\beta=\text{ReLU}(\text{BN}(\text{Conv}(F_{g}))),\$ (12)

### v3

# Breaking Down Monocular Ambiguity: Exploiting Temporal Evolution for 3D Lane Detection

@@ Abstract
Monocular 3D lane detection aims to estimate the 3D position of lanes from frontal-view (FV) images. However, existing methods are fundamentally constrained by the inherent ambiguity of single-frame input, which leads to inaccurate geometric predictions and poor lane integrity, especially for distant lanes. To overcome this, we propose to unlock the rich information embedded in the temporal evolution of the scene as the vehicle moves. Our proposed Geometry-aware Temporal Aggregation Network (GTA-Net) systematically leverages the temporal information from complementary perspectives. First, Temporal Geometry Enhancement Module (TGEM) learns geometric consistency across consecutive frames, effectively recovering depth information from motion to build a reliable 3D scene representation. Second, to enhance lane integrity, Temporal Instance-aware Query Generation (TIQG) module aggregates instance cues from past and present frames. Crucially, for lanes that are ambiguous in the current view, TIQG innovatively synthesizes a pseudo future perspective to generate queries that reveal lanes which would otherwise be missed. The experiments demonstrate that GTA-Net achieves new SoTA results, significantly outperforming existing monocular 3D lane detection solutions.
@@
Figure 1: Comparison with Monocular 3D Lane Detection Approaches. (a) Existing methods [huang2023anchor3dlane, bai2023curveformer, guo2020gen, chen2022persformer, garnett20193d] that rely on a single frame as input are hindered by inaccuracies in 3D geometry perception and difficulties in maintaining lane integrity. (b) In contrast, our approach, which incorporates multiple frames along with synthetic future frames, facilitates improved 3D geometric perception and enables integral lane detection within the scene.
@@ 1 Introduction
Monocular 3D lane detection, the task of estimating the precise 3D coordinates of lane lines from a single camera, is a cornerstone of modern autonomous driving systems [ma2024monocular]. Accurate 3D lane perception is crucial for safe lane keeping [wang2005lane], as well as for downstream tasks like high-definition map construction [li2022hdmapnet, liu2023vectormapnet] and trajectory planning [williams2022trajectory]. While cost-effective and rich in texture, monocular vision presents a formidable challenge: how to recover 3D structure from a 2D image.
Existing methods attempt to bridge this 2D-to-3D gap through various strategies, such as two-stage projection with depth estimation [yan2022once], or by transforming features into a bird’s-eye-view (BEV) representation [chen2022persformer, garnett20193d]. However, these approaches are built on fragile assumptions, like the reliance on highly accurate depth predictions or a flat-ground world, which often fail in complex real-world scenarios. More recent methods using 3D anchors or queries [huang2023anchor3dlane, luo2023latr] have shown promise, yet they still struggle with the fundamental ambiguity inherent in a single frame. This ambiguity manifests in two critical failures, as illustrated in Figure 1 (a): 1) inaccurate geometry, where predicted lanes deviate from their true 3D positions, and 2) broken lane integrity, where lanes, especially at a distance, are partially detected or missed entirely.
We argue that the key to overcoming these limitations lies not within the single frame, but in the temporal evolution of the scene. As a vehicle moves, the interplay between consecutive frames contains powerful cues to resolve the aforementioned ambiguities. In this paper, we propose the Geometry-aware Temporal Aggregation Network (GTA-Net), a novel framework designed to systematically exploit this temporal information from complementary perspectives, as shown in Figure 1 (b).
First, to address the challenge of inaccurate geometry, we ask: how can the past inform the present’s geometry? The relative motion between frames directly encodes the spatial relationships among scene elements. Inspired by this, we introduce the Temporal Geometry Enhancement Module (TGEM). By explicitly modeling the geometric consistency between past and present frames, TGEM learns to perceive depth-aware features, effectively correcting the geometric representation of the scene and leading to more accurate lane positioning.
Second, to tackle the problem of broken lane integrity, we ask: how can temporal context help us see ambiguous lanes? A lane that is faint or occluded now may have been clear in the past, or will become clear in the future. We therefore propose the Temporal Instance-aware Query Generation (TIQG) module. TIQG generates lane queries by aggregating instance information from both past and present views. More importantly, it addresses the challenge of detecting distant lanes with a novel strategy. As shown in Figure 4, a distant, undetectable lane becomes visible when the viewpoint moves closer. Since we cannot access real future frames, TIQG innovatively synthesizes a pseudo-future frame by digitally zooming into the region of interest. This allows us to generate effective queries for distant regions, dramatically improving lane integrity.
By seamlessly integrating these two components, GTA-Net leverages the temporal evolution to build a comprehensive understanding of the driving scene. Our main contributions are as follows:
@@
  • We propose the Geometry-aware Temporal Aggregation Network (GTA-Net), a novel architecture that systematically leverages temporal information to overcome the limitations of monocular 3D lane detection.
@@
  • We introduce the Temporal Geometry Enhancement Module (TGEM), which exploits geometric consistency across frames to learn depth-aware features for accurate 3D geometry perception.
  • We present the Temporal Instance-aware Query Generation (TIQG) module, which enhances lane integrity by aggregating temporal instance cues.
  • Extensive experiments show that GTA-Net achieves new state-of-the-art performance on the OpenLane benchmark, demonstrating its effectiveness and robustness across diverse and challenging scenarios.
@@ 2 Related Work
2D lane detection, a critical aspect of autonomous driving, strives to precisely determine the shapes and positions of lanes within images [tang2021review, xu2020curvelane]. With the development of deep learning [kirillov2023segment, yan2025olidm, yan2025drivingsphere, yan2025rlgf], deep neural network-based 2D lane detection has achieved remarkable strides forward [jin2023recursive, qin2022ultra, wang2022keypoint, zheng2022clrnet, han2022laneformer]. According to the formulation of 2D lane detection, existing solutions can be divided into four categories: segmentation-based, anchor-based, keypoint-based and curve-based methods [huang2023anchor3dlane].
Segmentation-based approaches [ko2021key, li2021abssnet, zheng2021resa] employ segmentation tasks as intermediaries for executing lane detection. To be specific, these methods firstly complete pixel-wise classification to obtain 2D lane segmentation masks. Next, the segmentation masks are processed by some carefully-design strategies to construct a set of lanes. Anchor-based approaches primarily utilize line-like anchors for regressing offsets relative to ground truth lanes [jin2022eigenlanes, liu2021condlanenet, qin2020ultra, qin2022ultra, tabelini2021keep, yoo2020end]. To enhance the efficacy of line-shape anchors, row-based anchors have been devised for capitalizing on the natural shape of lanes [yoo2020end]. By using row-based anchors, 2D lane detection task is converted into a row-wise classification problem. Furthermore, these row-based and column-based anchors are combined to reduce the occurrence of localization errors in side lanes [qin2022ultra]. Keypoint-based approaches offer a flexible framework for modeling lane positions [ko2021key, qu2021focus, wang2022keypoint]. At the outset, these methodologies embark on predicting the 2D coordinates of pivotal points along the lanes [ko2021key]. Following this initial step, a diverse array of strategies is employed to effectively group these key points into distinct lanes [wang2022keypoint]. Curve-based approaches leverage a range of curve formulations to model lanes [feng2022rethinking, liu2021end, tabelini2021polylanenet]. Specifically, this type of method aims to estimate the parameters of curves to obtain 2D lane detection results. BézierLaneNet [feng2022rethinking] leverages the parametric Bézier curve for its computational simplicity, stability, and extensive flexibility in transformations. Additionally, the authors introduce a novel feature flip fusion module, which takes into account the symmetrical characteristics of lanes.
Figure 2: Illustration of the Overall Architecture of the Proposed GTA-Net. The input images are first fed into a backbone network to extract 2D perspective features. Next, TGEM is applied to enhance the geometric perception. Simultaneously, TIQG is used to generate lane queries. The lane queries and features are sent to the decoder to obtain 3D lane predictions.
@@
The objective of 3D lane detection is to identify lanes within 3D space and accurately estimate their coordinates [ma2024monocular, zheng2024pvalane, li2023grouplane]. In contrast to 2D lane detection, 3D lane detection methods need to perceive the geometric information of scenes. Some methodologies try to enhance perception accuracy by leveraging both lidar and camera data [luo2023dv, luo2022m]. However, the acquisition and annotation expenses associated with multi-modal data are prohibitively high, constraining the practical implementation of such methodologies. Hence, monocular 3D lane detection garners increased interest due to its provision of a cost-effective alternative to multi-modal methodologies [ma2024monocular].
Monocular 3D lane detection is challenging due to the absence of depth information. To address this issue, SALAD [yan2022once] incorporates a depth prediction branch to estimate depth, which is used to project 2D lanes to 3D space. However, employing depth estimation in the process of detecting lanes may introduce cumulative errors and further exhibits inferior performance. IPM-based approaches have become increasingly prominent owing to their advantageous geometric properties for lane depiction in BEV perspective [ma2024monocular]. 3D-LaneNet [garnett20193d] utilizes IPM for the conversion of FV features into BEV features, enabling bidirectional information flow between image-view and top-view representations. Next, these features are leveraged for learning the anchor offsets of lanes in the BEV space. Gen-LaneNet [guo2020gen] introduces a novel lane anchor representation guided by geometry in a newly defined coordinate system, which employs a specific geometric transformation to directly compute real 3D lane points from the network output. It is essential to emphasize that aligning the lane points with the underlying top-view features in the newly defined coordinate frame is crucial for establishing a generalized method capable of effectively handling diverse scenes.
To overcome IPM’s reliance on the flat-ground assumption [chen2023efficient], recent methods operate directly in 3D. Anchor3DLane [huang2023anchor3dlane], for instance, forgoes the BEV representation by regressing 3D anchors directly, which are then projected onto the image to extract features. LATR [luo2023latr] uses a transformer-based architecture for monocular 3D lane detection. Specifically, the authors design a lane-aware query generator that produces queries derived from 2D lane segmentation. Besides, dynamic 3D ground PE generator is presented to embed 3D information during query updating.
@@ 3 Method
Our proposed Geometry-aware Temporal Aggregation Network (GTA-Net) leverages temporal evolution to address the core challenges of monocular 3D lane detection. This section details its architecture and key components. We first provide an overview of the entire GTA-Net framework. Then, we elaborate on our two core contributions: the Temporal Geometry Enhancement Module (TGEM) for better geometry perception, and the Temporal Instance-aware Query Generation (TIQG) for enhancing lane integrity. Finally, we describe the training objective and loss functions.
@@ 3.1 Overall Architecture
$L=(P,C),\quad P={(x_{i},y_{i},z_{i})}_{i=1}^{N},\\$ (1)
where $L$ denotes a 3D lane, $C$ represents the corresponding lane category of $L$, $P$ is composed of $N$ 3D points with $N$ being a predefined parameter. It is noted that the sequence $\{y_{1},y^{2},...,y^{N}\}$ is often defined as a predetermined longitudinal coordinate [garnett20193d, guo2020gen, chen2022persformer]. Based on the above descriptions about lane representation, the 3D lanes in an input image $I_{t}$ can be represented as $G=\{L^{1},L^{2},...,L^{M}\}$, where $G$ and $M$ represent the collection and the number of 3D lanes, respectively.
@@
$F_{t}=\text{Backbone}(I_{t}),\quad F_{t-n}=\text{Backbone}(I_{t-n}),\\$ (2)
@@
$Q=\text{TIQG}(F_{t-n},F_{t}),\\$ (4)
where $Q$ represents the generated instance-aware queries, and $\text{TIQG}(\cdot)$ denotes the transformation performed by the TIQG module. Next, the instance-aware queries $Q$ and geometry-aware features $F_{ge}$ are fed into a deformable attention [xia2022vision, zhu2020deformable] based decoder to iteratively update queries:
$Q_{u}=\text{Decoder}(Q,F_{ge}),\\$ (5)
where $Q_{u}$ and $\text{Decoder}(\cdot)$ denote the updated queries and the transformation of the above deformable attention-based decoder, respectively. Finally, the updated queries $Q_{u}$ are fed into the 3D lane detection head to output the predictions of 3D lanes in the input image $I_{t}$:
$\hat{G}=\text{Head}(Q_{u}),\\$ (6)
where $\hat{G}$ indicates the estimated 3D lanes, and $\text{Head}(\cdot)$ denotes 3D lane detection head. For the losses, we follow the setting of [luo2023latr].
@@ 3.2 Temporal Geometry Enhancement Module
The inherent absence of depth cues in monocular input images poses a significant challenge for monocular 3D lane detection solutions. Without depth information, accurately understanding the complex geometric details of driving scenes becomes difficult. Drawing inspiration from prior research [watson2021temporal, yang2020cost] in monocular depth estimation, this study aims to alleviate this limitation by utilizing the geometric relationship between consecutive frames. To this end, we propose Temporal Geometry Enhancement Module (TGEM), which is designed to delve into the geometric consistency hidden in successive frames. The structure of TGEM is described by figure 2. In detail, TGEM consists of three main stages: cost volume construction, geometric feature leaning, geometry-aware feature enhancement.
Cost volume construction. To uncover the potential geometric information within successive monocular images, we first construct a cost volume in the initial stage of TGEM. This is because the cost volume is capable of capturing the geometric relationship between adjacent frames [gu2020cascade].
@@
$E_{t}=\text{Repeat}(F_{t}),\quad E_{t-n}=\text{Repeat}(F_{t-n}),\\$ (7)
where $E_{t}$ and $E_{t-n}$ denote the expanded features, and $\text{Repeat}(\cdot)$ represents the feature expanding operation. Next, the expanded features $E_{t-n}$ in image coordinate system of tnt-n frame is projected onto image coordinate system of $t$ frame by using camera intrinsics, extrinsics and lidar poses:
$\hat{E_{t}}=\text{Warp}(F_{t-n}),\\$ (8)
where $\hat{E_{t}}$ denotes the warped features in image coordinate system of $t$ frame, and $\text{Warp}(\cdot)$ indicates the projection from image coordinate system of tnt-n frame to image coordinate system of $t$ frame. In the final step, the matching cost between $\hat{E_{t}}$ and $E_{t}$ is computed by using $l_{1}$ norm:
$C=\lVert\hat{E_{t}}-E_{t}\rVert_{1},\\$ (9)
@@
$F_{g}=\text{ReLU}(\text{BN}(\text{Conv}(C))),\\$ (10)
@@
Figure 3: An Overview of the Proposed Temporal Instance-aware Query Generation (TIQG) Module. TIQG strategically integrates temporal cues into the query generation process, thereby enhancing its capacity to exploit temporal instance information for promoting lane integrity.
@@
$\alpha=\text{Sigmoid}(\text{BN}(\text{Conv}(F_{g}))),\\$ (11)
$\beta=\text{ReLU}(\text{BN}(\text{Conv}(F_{g}))),\\$ (12)